OpenStudy (anonymous):

If f(x) =x-3 and g(x) =2x-4, find (f+g)(x)

OpenStudy (anonymous):

also Find the zero of f(x) =5x-2

OpenStudy (anonymous):

add f(x) with g(x)

OpenStudy (anonymous):

i just combine them?

OpenStudy (anonymous):

yes,(f+g)(x)=f(x)+g(x)

OpenStudy (anonymous):

alright do you know the other one?

OpenStudy (anonymous):

5x-2=0 then solve it

OpenStudy (anonymous):

okk so the first one would be (f+g)(x) =3x-7? and the second would be 5x-2?

OpenStudy (anonymous):

5x-2=0 5x=2 x=2/5

OpenStudy (anonymous):

but is the first one right then?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

alright thank you!